OpenStudy (kcnkinoie):

Write the equation of the line that passes through (2, 4) and has a slope of –1 in point-slope form. x – 4 = –1(y – 2) x – 2 = –1(y – 4) y – 2 = –1(x – 4) y – 4 = –1(x – 2)

OpenStudy (wckeller):

Point slope form:\[y - y_1 =m(x-x_1)\] Plug in your point, (2,4) and the slope m=-1 \[y-4=-1(x-2)\] The correct answer is the last option on the list you provided.